# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ A.12 Tiny-Imagenet results
In an effort to demonstrate the potential of LAP further, we evaluate our method on Tiny-Imagenet [32], a subset of Imagenet that contains $100000$ images from $200$ classes. In this experiment, we wanted to test the use of LAP on larger images with larger numbers of sources. Firstly, we load the training and testing sets and randomly split the training set into $100$ sources uniformly. We then choose $40$ of the sources to be $100\%$ unreliable, and introduce noise through random labelling. This dataset is then split in the ratio $9:1$ to produce a validation set. For this experiment, we train a ResNet 50 architecture using the baseline training method provided for Imagenet in Pytorch  [^11] https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/. This trains the model for $90$ epochs with a batch size of $256$ using stochastic gradient descent with an initial learning rate of $0.1$, momentum of $0.9$, and weight decay of $0.0001$, as well as a learning rate scheduler that multiplies the learning rate by $0.1$ every $30$ epochs. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources.
The results of this are available in Table 7 and again demonstrates the expected performance improvement when using LAP for datasets with sources of unknown noise, and the maintenance of performance when data is non-noisy.
@@
A.13 Imagenet results with multiple noise types
Model Types
Noise Type Standard LAP (Ours)
Input and Label Noise 68.05 ± 0.26 70.61 ± 0.26
We additionally present our results on Imagenet, which allows us to test our method on a large scale dataset. Within this experiment, as in that presented in Figure 3(a), we allow for both input and label noise. In Figure 3(a) all noisy sources contain two types of noise, however in this experiment we allow for different sources to contain different types of noise. We further observe, in Table 8, that using LAP improves model performance here, since it achieves a greater maximum top-5 test accuracy as expected.
For this experiment, we train a ResNet 50 architecture using the baseline training method provided for Imagenet in Pytorch11. This trains the model for $90$ epochs with a batch size of $256$ using stochastic gradient descent with an initial learning rate of $0.1$, momentum of $0.9$, and weight decay of $0.0001$, as well as a learning rate scheduler that multiplies the learning rate by $0.1$ every $30$ epochs. We split the Imagenet data into $10$ sources, of which $5$ are chosen to be noisy ($2$ containing input noise and $3$ containing label noise). Images are loaded as 8-bit integer arrays and interpolated to $(64,64)$ before input noise is synthesised by randomly adding uniform integers from [-64, 64]. They are then transformed to 32-bit floats and normalised using the mean and standard deviation available on the same Pytorch training script. Label noise is added by randomly replacing labels in noisy sources. This dataset is then split in the ratio $9:1$ to produce a validation set for the analysis in Appendix A.13.1. A version of the ResNet 50 model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. These experiments are repeated $5$ times for each setting.
A.13.1 Late training test accuracy
@@
The noisy data literature often presents model performance results on the last $x$ epochs, since in reality (as we don’t have access to clean test labels) there is no way of knowing when to stop training and reduce overfitting to noisy data. In our work, in an effort to be most fair to the standard training baseline, we present the maximum performance over all of training to ensure that we are not presenting results in which the standard training baseline has significantly overfit to the noisy data. However, in Table 9 we also present the average test accuracy over the last $10$ epochs as is often reported in the literature. These results show the significant improvement that using LAP can have on model training when it is unknown if the standard training method is overfitting to noisy data. Similarly, Figure 5 presents the training, validation, and testing performance at each epoch during the training of a neural network using the standard method and LAP. We can see that the training curves appear usual (Figure 5(a), with LAP’s training curve representing the tempered loss), and that the validation accuracy is greater for standard training (since the validation set contains noisy labels), but that when tested on the clean labels (Figure 5(c)), it is clear that using LAP enables considerably more robustness to noisy data. Early stopping could be used here to achieve the maximally achieving models (since the epoch of maximum accuracy on the test set is the same as the validation set), which is why we find it more informative to report the maximum test accuracy as presented in all other experiments.
(a) Training cross entropy loss.
(b) Validation top-5 accuracy.
(c) Test top-5 accuracy.
Figure 5: Performance on Imagenet. In 5(a) we show the cross-entropy loss on the training data at each epoch where the loss using LAP represents the tempered cross-entropy loss. In 5(b) we present the validation top-5 accuracy at each epoch. In 5(c) we present the test top-5 accuracy at each epoch to demonstrate the fitting to the noise that occurs when not using LAP during late training. Here, $5$ out of a total of $10$ sources are $100\%$ noisy, with $2$ containing input noise, and $3$ containing label noise. The two steps in performance occur at the points at which we scale the learning rate using a scheduler (epoch $30$ and $60$). The lines and error bands represent the mean and standard deviation over the $5$ repeats. The vertical black dashed line represents the epoch at which LAP achieved maximum top-5 accuracy on the test set.
A.14 California Housing: Regression results
Table 10: Standard and LAP results on a regression task. Mean ± standard deviation of minimum mean squared error of standard training and LAP over $5$ repeats on the California Housing dataset with different types of noise. Here, $4$ out of $10$ sources are 100% noisy. Note that “Random Label" is the only noise type with a significant difference between methods. Here, smaller is better.
@@ A.14 California Housing: Regression results
In Table 10 we present the results of using LAP and standard training on a regression dataset in which we have $10$ total sources and $4$ that are $100\%$ noisy, with labels replaced with uniform noise (sampled between the minimum and maximum label value in the training set).
To test our method against standard training, we evaluate both on a dataset of $20640$ samples of house values in California districts, with the goal to predict the median house value from the U.S census data for that region – a regression task. We randomly split the data into training and testing with an $8:2$ ratio for each of the $5$ runs we performed over each noise type. We train a multilayer perceptron as described in Section A.7 for $200$ epochs with a batch size of $256$, leaning rate of $0.001$ using stochastic gradient descent with momentum of $0.9$ and weight decay of $0.0001$. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. This was repeated $5$ times for each setting.
Table 10 shows that LAP improves the test set performance significantly over the baseline for random labelling noise and maintains performance when no noise is present. This is as expected, and again illustrates its effectiveness, but this time on a regression task.

### v4

@@ A.17 Tiny-Imagenet results
In an effort to demonstrate the potential of LAP further, we evaluated our method on Tiny-Imagenet [34], a subset of Imagenet that contains $100000$ images from $200$ classes. In this experiment, we wanted to test the use of LAP on larger images with larger numbers of sources. Firstly, we load the training and testing sets and randomly split the training set into $100$ sources uniformly. We then choose $40$ of the sources to be $100\%$ unreliable and introduce noise through random labelling. The dataset is then split into the ratio $9:1$ to produce a validation set. For this experiment, we train a ResNet 50 architecture using the baseline training method provided for Imagenet in Pytorch  [^13] https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/. This trains the model for $90$ epochs with a batch size of $256$ using stochastic gradient descent with an initial learning rate of $0.1$, momentum of $0.9$, and weight decay of $0.0001$, as well as a learning rate scheduler that multiplies the learning rate by $0.1$ every $30$ epochs. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources.
The results of this are available in Table 11 and again demonstrate the expected performance improvement when using LAP for datasets with sources of unknown noise, and the maintenance of performance when data is non-noisy.
@@
A.18 Imagenet late training test accuracy
@@
The noisy data literature often presents model performance results on the last $x$ epochs, since in reality (as we do not have access to clean test labels) there is no way of knowing when to stop training and reduce overfitting to noisy data. In our work, in an effort to be most fair to the standard training baseline, we present the maximum performance over all of training to ensure that we are not presenting results in which the standard training baseline has significantly overfit to the noisy data.
However, in Table 12 we also present the average test accuracy over the last $10$ epochs, as often reported in the literature. These results show the significant improvement that the use of LAP can have on model training when it is unknown if the standard training method is overfitting to noisy data. Similarly, Figure 9 presents the training, validation, and testing performance at each epoch during the training of a neural network using the standard method and LAP. We can see that the training curves appear as usual (Figure 9(a), with LAP’s training curve representing the tempered loss) and that the validation accuracy is greater for standard training (since the validation set contains noisy labels), but that when tested on the clean labels (Figure 9(c)), it is clear that using LAP enables considerably more robustness to noisy data. Early stopping could be used here to achieve the maximally achieving models (since the epoch of maximum accuracy on the test set is the same as the validation set), which is why we find it more informative to report the maximum test accuracy as presented in all other experiments.
A.19 California Housing: Regression results
Table 13: Standard and LAP results on a regression task. Mean ± standard deviation of minimum mean squared error of standard training and LAP over $5$ repeats on the California Housing dataset with different types of noise. Here, $4$ out of $10$ sources are 100% noisy. Note that “Random Label” is the only noise type with a significant difference between methods. Here, smaller is better.
@@ A.19 California Housing: Regression results
In Table 13 we present the results of using LAP and standard training on a regression dataset in which we have $10$ total sources and $4$ that are $100\%$ noisy, with labels replaced with uniform noise (sampled between the minimum and maximum value of the label in the training set).
To test our method against standard training, we evaluate both on a dataset of $20640$ samples of house values in California districts, with the goal to predict the median house value from the U.S census data for that region – a regression task. We randomly divide the data into training and testing with a $8:2$ ratio for each of the $5$ runs we performed over each noise type. We train a multilayer perceptron as described in Section A.9 for $200$ epochs with a batch size of $256$, leaning rate of $0.001$ using stochastic gradient descent with momentum of $0.9$ and weight decay of $0.0001$. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. This was repeated $5$ times for each setting.
Table 13 shows that LAP improves the performance on the test set significantly over the baseline for random labelling noise and maintains the performance when no noise is present. This is as expected and again illustrates our method’s effectiveness, but this time on a regression task.
@@
A.20 Additional Compute Example
In this section, we briefly describe the additional cost of applying LAP with a specific example.
In practise using our implementation, with an MLP with hidden sizes $20-2000-2000-5$, and batches of $512$ samples (with 20 features, 5 classes, and data generated from 10 sources) the forward-backward pass through the model (without source reweighting) takes 15300 $\mu$ s ± 164 $\mu$ s and the reweighting of losses takes 469 $\mu$ s ± 10.8 $\mu$ s, increasing the time by  3%. With 50 and 100 sources, the reweighting of losses takes 1000 $\mu$ s ± 11.3 $\mu$ s and 1990 $\mu$ s ± 49.2 $\mu$ s respectively.
Using the same set-up, with batches of 2048 in size and 10 sources, the forward-backward pass (without reweighting) takes 41000 $\mu$ s ± 689 $\mu$ s whilst the loss reweighting takes 467 $\mu$ s ± 8.51 $\mu$ s (no significant change as the batch size increases), increasing the time to compute the loss and gradients on a batch by  1%.
We believe that this is a fair trade-off for the potential improvement in performance.